Bill and Jane are mowing a lawn. If bill mowed 1/8 of it and Jane mowed 1/4, how much still remains?

Answer by ewatrrr(24785)


Hi
Bill and Jane are mowing a lawn.
If bill mowed 1/8 of it and Jane mowed 1/4
how much still remains? 1 - 1/8 -1/4 = 1-1/8 -2/8 = 1- 3/8 = 5/8 remains